In 3 months Ann needs to repay $200.How much did she borrow today if the bank is charging 9% per annum on the loan?
1
Expert's answer
2018-04-04T04:50:29-0400
Solution:
Let Ann did borrow today x dollars. Then she needs to repay in 1 month:
x+x*1 months *9% /12 months = x*(1+0,0075) = x * 1,0075;
Also, she needs to repay in 2 month:
x * 1,0075*1,0075.
And she needs to repay in 3 month:
x * 1,0075*1,0075*1,0075 = 200 dollars.
It means:
x = 200/ 1,0075 3 ;
x = 195,57 dollars.
Answer:
Ann did borrow today 195,57 dollars.